The volume of a box V, varies with some variable x as V(x) = x3-12x2+44x-48 cubic metres. If (x - a) metre is the measurement of one side of the box, then find the value for a


1
Expert's answer
2021-10-06T17:51:54-0400
"V(x)=x^3 - 12x ^2 + 44x -48""x^3 - 12x ^2 + 44x -48 =x^2(x-2)-10x(x-2)+24(x-2)""=x^2(x-2)-10x(x-2)+24(x-2)""=(x-2)(x^2-10x+24)""=(x-2)(x-4)(x-6)""a_1=2, a_2=4, a_3=6""a\in\{2,4,6\}"